Calculate the mass defect and nuclear binding energy per nucleon of each of the nuclides
a.O-16(atomic mass=15.994915 amu)
b.Ni-58(atomic mass=57.935346 amu)
c.Xe-129(atomic mass=15.994915 amu)
1
Expert's answer
2020-05-05T18:40:55-0400

a)

"Z=8, N=A-Z=16-8=8"


"\Delta m=8(1.00783)+8(1.00866)-15.994915\\=0.137005\ amu"

"E=(0.137005)\frac{931.5}{16}=7.976\frac{MeV}{nucleon}"

b)


"Z=28, N=A-Z=58-28=30"

"\Delta m=28(1.00783)+30(1.00866)-57.935346 \\=0.534694\ amu"

"E=(0.534694)\frac{931.5}{58}=8.732\frac{MeV}{nucleon}"

c)


"Z=54, N=A-Z=129-54=75"



"\Delta m=54(1.00783)+75(1.00866)-128.904780\\=1.16754\ amu"


"E=(1.16754)\frac{931.5}{129}=8.431\frac{MeV}{nucleon}"